A woman is reported to have fallen 121 ft from a building, landing on a metal ventilator box, which she crushed to a depth of 26 in. She suffered only minor injuries. (a) Neglecting air resistance, calculate the speed of the woman just before she collided with the ventilator box. ft/s (b) Calculate her average acceleration while in contact with the box (c) [Not Graded] Calculate the time it took to crush the box.
Answer #1

Use of equation of motion...Solution supfose the speed of te women v Just belore ceuidig the buxs ヲ 88.9 sec Her overage accclovotion while in ContereG with ox woment speed oteduced to zexo after descendi aheight o R in So 12 (a= lt/s-l.oeroS ietotke. ]824-30 (C) Time et took to crush the box÷ づ 0 = 88.91-1839-30Kt. づ -t 88.91 1824.3o t= 0.048 sec